OpenStudy (anonymous):

what is 124% writen as a fraction?

OpenStudy (anonymous):

please help me i dont under stand

OpenStudy (anonymous):

31/250

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

google search

OpenStudy (anonymous):

i need to understand how to do the problem not just what the answer it :|

OpenStudy (anonymous):

@realmadridCR7 thst answer is wrong

OpenStudy (anonymous):

*that

OpenStudy (anonymous):

ok can u show me what it is

OpenStudy (anonymous):

do you know how to turn 124% into a decimal ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

take 124 and divide it by 100 and what do you get ?

OpenStudy (anonymous):

1.24

OpenStudy (anonymous):

sure. Since its 124%, it simply means \[\left(\begin{matrix}124 \\ 100\end{matrix}\right)\] therefore in its lowest term, you just reduce it so that will be \[\left(\begin{matrix}31 \\ 25\end{matrix}\right)\] ...In order to arrive at that answer , you simply divide both the numerator and denominator by the lowest common factor, 4 so.... 124 divided by 4 is 31 and 100 divided by 4 is 25. There the answer is \[\left(\begin{matrix}31 \\ 25\end{matrix}\right)\]

OpenStudy (anonymous):

wow, thanks!

OpenStudy (anonymous):

I usually turn it into a decimal first then into a fraction. Not everybody does it that way

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

oh yah i think it would be easyer for me if i just learned how to just turn a percent into a fraction

OpenStudy (anonymous):

whatever is easier for you is what you need to stick with :)

OpenStudy (anonymous):

yep any number with the % sign means over 100. Always remeber that.because its much easier that way :)

OpenStudy (anonymous):

:) ok.